# Mapping the Podcast Ecosystem with the Structured Podcast Research Corpus

### v1

@@ 4 Mapping The Podcast Ecosystem
Although there are many studies of podcast content Lindgren (2016); Drew (2017); Little et al. (2020); Morris and Hoyt (2021), advertising Moe (2022); Brinson and Lemon (2023); Bezbaruah and Brahmbhatt (2023), audiences Pew Research Center (2023c); Werner et al. (2020), and impacts Wirtschafter (2023); Pew Research Center (2023d), all of these have been carried out at a small scale, often relying on curated samples and/or manual annotation. To provide the first comprehensive documentation of the overall podcast ecosystem, we begin by characterizing its content and network structure. In particular, we focus on three questions: 1) What content appears on podcasts? 2) To what extent is this content organized into coherent communities? and 3) How are podcasts connected within and across categories?
Our question regarding podcast content is motivated by the fact that information in podcasts has had both helpful and pernicious effects on society. In the medical domain, for example, podcasts have served as a supportive tool for education and support Billman et al. (2024); Hurst (2019) as well as source of dangerous misinformation, particularly during the height of the COVID-19 pandemic Wirtschafter (2023). To facilitate research into these areas it is important to investigate what topics podcast creators discuss, and how concentrated these topics are within categories.
To complement our topical analysis, we capture the social network of this ecosystem, using guests as key links between shows. Prior work has found that edges in social networks facilitate the diffusion of information Bakshy et al. (2012). Furthermore, guest invitations reflect the preferences of podcast hosts and their audiences. Thus, podcasts linked through shared guests may play a role in forming communities with a shared identity or set of common knowledge. Prior work has demonstrated this at a very small scale Stocking and Odabas (2024), and we extend this work to provide the macro-level analysis crucial to understanding the overall structure of this ecosystem.
@@ Methods
Podcasts have self-described high-level category labels (e.g., News, Sports, Society). To understand the interaction between content and category, we augment these labels to directly model the content in episodes. We fit an LDA topic model on the first 1000 words of each podcast episode transcript with 200 topics, using Mallet (Blei et al., 2003; McCallum, 2002). Each episode is then represented as its topic distribution, allowing us to compare episodes in terms of topical similarity.
@@
Guests are identified as any two-word named entity labeled as Guest by our model, as described in §3.4. To minimize the effect of false positives, we exclude guests that have names among the top 50% of most frequent named entities (e.g., John Smith) as well as all names classified as a Host in a different episode of the same podcast. We then project this bipartite network to a one-mode network where podcasts are connected if they share one or more guests in common, which we refer to as the podcast-guest network. The resulting network has 10,480 vertices (podcasts) and 26,589 edges (connections based on co-appearing guests). For details, see Appendix F.
Topical Distribution
Visualizing podcasts by their content similarity reveals coherent topical communities, as well as cross-cutting themes that span multiple categories, as seen in Figure 1. Many topical clusters fall neatly within their self-ascribed category labels, indicating that these categories are a meaningful descriptor of the type of content within podcasts. This is particularly true for categories such as Sports, Religion, and Business, each of which have coherent thematic subclusters such as Baseball, Wrestling, Real Estate, Bitcoin, Judaism and Islam. Other broad topics appear to be further fragmented, such as Christianity, for which a number of topics appear to correspond to different parts of the liturgy or themes within Christianity.777Religion turns out to be the most common podcast category in our dataset, with the vast majority of episodes in this category taking the form of recorded Christian sermons.
By contrast, however, there are a number of topics that appeal to creators from multiple categories or supersede category boundaries entirely. Two such clusters focus on spirituality and self-improvement, which include content from episodes in the religion, society, and business domains and the Business, Religion, and Health categories respectively. This is also the case for topics related to COVID-19, as well as racial justice, which is explored in-depth in §5. The existence of such communities suggests that existing category labels do not fully describe the meaningful groupings in podcast content, and that these are the areas in which we might expect to see the most extensive cross-fertilization and exchange of ideas.
A different, but suggestive measure of categorical homogeneity is provided by vocal characteristics. As expected, podcasts in the Kids category (aimed at children) have the highest average pitch. By contrast, we find the lowest average pitch in Science, Leisure, and Technology categories, a pattern which may reflect a relative lack of gender diversity (see Appendix Figure 8).
@@
Guests play an important role in structuring the communities within podcasts. Guest appearances are extremely common, with Stocking and Odabas (2024) finding over 7,000 guest appearances on a set of 434 top-ranked podcasts in 2022. High-profile guests may appear frequently on multiple podcasts to promote their brand or ideology Bratcher and Cabosky (2024) facilitating the exchange or even diffusion of ideas. As a result, guest appearances are often a strong indicator about what type of content hosts prefer to feature and what will appeal their audience.
We find that the association between guest appearances and topical content varies by the high-level category of podcast, as shown by the podcast-guest network in Figure 2. Here, we visualize the network with nodes colored by category, positioned by the force-directed Yifan Hu algorithm, and sized proportionate to their edge degree Hu (2005). The figure shows that some categories such as Business, News, and Sports categories correspond to relatively distinct modules—i.e., podcasts within the same category share a distinct pool of guests that do not interact much with podcasts from different categories. Despite being the largest categories, Religion and Society do not form large clusters in our network. As shown in supplemental Figure 6, both the Religion and Society categories appear to invite guests substantially less than other large podcast categories, and thus provide fewer opportunities for cross-fertilization of ideas. This suggests that communities may be less well connected for these categories, though a longitudinal sample may help identify such structure.
@@
To quantify this claim, we calculate modularity for the partitions of our network corresponding to different categories to measure their intraconnectedness. This measures the extent to which membership in a category increases the probability of two nodes sharing a common guest relative to a graph with randomized edges (see Appendix G for details). The high modularity of Business and Sports suggests that podcasts within these categories are particularly likely to share guests relative to podcasts in other categories (Table 1).
@@
Overall, we find evidence that categories provide meaningful structure to the podcast ecosystem. With respect to content, we find many distinct fandoms represented (e.g., Wrestling, Bitcoin), with many such topics belong to a single category. However, a number of high-stakes topics (i.e. COVID-19, racial justice, and wellness) have cross-category discussion, a potential indicator that these domains may serve as a ground for discussion, debate, or the diffusion of ideas. With respect to guests, we find that business and sports form distinct communities, whereas two of our largest categories (religion, society) do not. Our results cannot by themselves inform how community structure influences discourse in the podcast sphere. However, they answer the most fundamental question for future work by telling us where this community structure exists.
@@ 5 Collective Attention in the Ecosystem
In news domains, media environments often respond rapidly to major events, setting the agenda for public discussion and debate King et al. (2017); Boydstun et al. (2014). Due to their ubiquity, podcasts may play an equally important role in agenda-setting. However, given their fragmentation, it’s not currently known how responsive the ecosystem is to current events. Here, we test for collective attention using a case study on George Floyd’s murder, measuring both the (1) temporal responsiveness and (2) overall diffusion associated with this event. George Floyd’s death was one of the most significant events of that year in terms of news coverage Cowart et al. (2022); Reny and Newman (2021), and, as it occurs in the middle of the SPoRC time frame, provides an ideal test case of the response of the podcast ecosystem to real world events.
@@ Results
Collective attention in podcasts responds rather quickly to the initial event, then decreases far slower over the following weeks (Figure 3). In particular, the percentages for both the George Floyd and Black Lives Matter topics rises to their peaks over the course of roughly 10 days, with a lag of approximately 4 days in-between them. This pattern of quick story development and slow story decay aligns with the phenomenon of a “media storm” in the news medium Boydstun et al. (2014). However, this case suggests that development of media storms in podcasts may operate at a slower pace, as prior computational work on news-media found that the average storm had peaked by day three and decayed by day ten Litterer et al. (2023). Given that there are substantially more podcasts than media outlets (in part due to ease of entry), future work is needed to understand how the two ecosystems differ in their responses.
@@
Our analyses confirm that 1) the podcast ecosystem’s response had similar temporal patterns to news-media, and that 2) discussion of George Floyd was widespread, yet variable, across communities. In the first case, our finding builds off of work on “media-storms” and builds toward future work into the underlying mechanisms through which collective attention in podcasts rapidly shift and slowly decay. With regard to the widespread discussion of George Floyd, our work is broadly supported by findings that political discussion is abundant across communities with both political and non-political labels Rajadesingan et al. (2021); Munson and Resnick (2011). This suggests that, as with other platforms, listeners may be incidentally exposed to political content from hosts whom they trust and feel connected to Pew Research Center (2023d, a); Fletcher and Nielsen (2018); Weeks and Lane (2020).
@@ 6 Discussion
In addition to NLP applications, our data and analyses enable deeper investigation into social science questions. Section 4 emphasized structure, identifying communities within podcasts. Future research could investigate the political and social properties of these communities, similarly to past work on the Reddit and News ecosystems Waller and Anderson (2021); Niculae et al. (2015). For example, our textual and prosodic information could be used to infer whether communities form along the lines of gender performance and identity in this male-dominated space.
In section 5, we illustrate collective attention dynamics in the podcast ecosystem through our study of racial justice discussions in 2020. Rather than focusing on a single event, future work may consider how this ecosystem responds to a number of politically salient topics, extending work on news-media to the podcast ecosystem Boydstun et al. (2014); Litterer et al. (2023). We found that discussion of George Floyd permeated even categories such as religion and business. Building on this finding, future work could model how information diffuses across communities and under what conditions this is likely to occur.
@@ 7 Conclusion
Podcasts are a widely distributed, unique, and impactful form of media, but comprehensive data for the computational study of this medium have been lacking. In this work, we present SPoRC, a dataset of over 1 million podcast transcripts alongside foundational analyses for the continued study of the podcast ecosystem. Our data is comprehensive, characterizing this medium in terms of its content, audio characteristics, speaker roles, and a number of other relevant factors such as category, hosting service, and publication date. Our structural analysis describes what content is discussed by which categories and demonstrates that podcast communities are formed through shared guests. Our temporal analysis finds that the ecosystem responded similarly to how the news-media respond to major events. Furthermore, discussion of George Floyd was widespread, reaching a diverse set of categories and a large fraction of total podcasts. Our data and analyses answer important questions about podcasts. More importantly, however, they open the door to the continued research of this impactful and understudied medium.
@@
Our data and analysis are limited in a number of ways. With regard to our dataset, there are certain podcasts that are not publicly available and therefore excluded from our data collection pipeline. One noteworthy example is the popular Joe Rogan podcast, which is available exclusively through Spotify. Moreover, despite being a large online repository, Podcast Index is not guaranteed to have comprehensive coverage of existing podcasts.
Our data processing pipeline involves additional limitations. Whisper models are predictive tools, meaning that they are inherently subject to mistakes and even hallucinations. These issues may be exacerbated in the context of non-majority accents, code-switching, or audio recorded with cheap or low quality equipment. In order to maximize data quality, we remove cases of potential hallucination, but this may disproportionately affect already marginalized groups. Furthermore, our diarization and role-annotation approaches are subject to error that is difficult to quantify and could propagate to downstream analyses in unforeseen ways. As noted, some hosts do not introduce themselves, hence we are unable to identify them, and we do not attempt to resolve names to speakers in the case of more than one host or guest.
While we provide a first large scale analysis of podcasts, our conclusions are limited. In our analyses, we use a subset of data that is dense but represents a narrow slice of time. Thus, our conclusions may not generalize outside this timeframe.
Further limitations with regard to our temporal analysis stem from our inability to pinpoint mechanistic factors underlying our observed time series. While prior work is suggestive of particular explanations for the patterns observed within and across categories, our data do not permit a definitive answer for “why” these patterns emerge.
@@
At the data-processing phase, one potential harm arises if a speaker’s actual speech is mistranscribed or misattributed. Such a mistake would be particularly problematic if this speech is offensive or reflects stereotypes or other biases. Furthermore, this type of error could be more likely for users with non-majority ways of speaking, which could exacerbate this harm for groups that are already marginalized. To mitigate this issue at the transcription level, we have undertaken thorough manual review to ensure that hallucinations are rare and that repeated text is filtered from our data. In the case of diarization, this issue is more difficult to mitigate. Recognizing that the potential for harm due to speaker misattribution varies widely across contexts, we encourage future users of our data to consider how this potential error could manifest when interpreting downstream results.
We further identify two potential harms due to the open release of our data. First, releasing our data could result in privacy concerns. Although we only process podcasts that are publicly available, the release of our processed data could expose a podcast’s content to an unforeseen or undesired audience. Creators may have expectations that their content will not be widely distributed, and the release of our data could violate this expectation. Furthermore, the enhanced search-ability afforded by podcast transcription could render a podcast’s content easier to access than was originally intended. A second and related harm is the usage of data for training of large language models. Some creators may not want their content being used to train deep-learning system, and the release of our data makes content more accessible for these purposes. This issue is particularly concerning when content is used to generate revenue without compensating original creators.

### v2

@@ 4 Mapping The Podcast Ecosystem
Although there are many studies of podcast content Lindgren (2016); Drew (2017); Little et al. (2020); Morris and Hoyt (2021), advertising Moe (2022); Brinson and Lemon (2023); Bezbaruah and Brahmbhatt (2023), audiences Whipple et al. (2023); Werner et al. (2020), and impacts Wirtschafter (2023); Pew Research Center (2023b), all of these have been carried out at a small scale, often relying on curated samples and/or manual annotation. To provide the first comprehensive documentation of the overall podcast ecosystem, we begin by characterizing its content and network structure. In particular, we focus on two questions: 1) What are the major content areas discussed on podcasts? and 2) To what extent are communities isolated or connected across different categories in terms of their topical coverage and common guests? Here, we distinguish between topical and networked communities, where topical communities are podcasts with similar content, and networked communities are podcasts linked through common guests.
To answer these questions, we consider both podcasts’ content and the social network formed by common guests. Our topical approach is motivated by the fact that information in podcasts has had both helpful and pernicious effects on society. In the medical domain, for example, podcasts have served as a tool for education and support Billman et al. (2024); Hurst (2019) as well as source of dangerous misinformation, particularly during the height of the COVID-19 pandemic Wirtschafter (2023). Our analysis here helps to reveal, for the first time at this scale, the overall prevalance of various topics of discussion on podcasts, both within and across categories.
To complement our topical analysis, we capture the social network of this ecosystem, using guests as key links between shows. Prior work has found that edges in social networks facilitate the diffusion of information Bakshy et al. (2012). Furthermore, guest invitations reflect the preferences of podcast hosts and their audiences. Thus, podcasts linked through shared guests may play a role in forming networked communities with a shared identity or set of common knowledge (Stocking and Odabas, 2024). In contemporaneous work, DeMets and Spiro (2025) have also looked at both the guest co-occurrence network, and the movement of guests between shows, using the PPPD corpus of around 120 mostly political podcasts. Whereas topics are the sole focus of our first research question, we use both topical information and this social network to answer the second.
@@ Methods
Podcasts have self-selected high-level category labels (e.g., News, Sports, Society), with most shows using well established primary categories. To understand the interaction between content and category, we augment these labels to directly model the content in episodes. We fit an LDA topic model on the first 1000 words of each podcast episode transcript with 200 topics, using Mallet (Blei et al., 2003; McCallum, 2002). Each episode is then represented as its topic distribution, allowing us to compare episodes in terms of topical similarity.
@@
Guests are identified as any two-word named entity labeled as Guest by our model, as described in §3.4. To minimize the effect of false positives, we exclude guests that have names among the top 50% of most frequent named entities (e.g., John Smith) as well as all names classified as a Host in a different episode of the same podcast. We then project this bipartite network to a one-mode network where podcasts are connected if they share one or more guests in common, which we refer to as the podcast-guest network. The resulting network has 10,480 vertices (podcasts) and 26,589 edges (connections based on co-appearing guests). For details, see Appendix G.
Topical Communities
To address the topical dimension of our research questions, we visualize podcasts with respect to their category label and content similarity, revealing both coherent topical communities as well as some that span multiple categories. Many topical communities fall neatly within their self-ascribed category labels, indicating that these categories are a meaningful descriptor of the type of content within podcasts. This is particularly true for categories such as Sports, Religion, and Business, each of which have coherent thematic sub-communities such as Baseball, Wrestling, Real Estate, Bitcoin, Judaism, and Islam.999Surprisingly, Religion turns out to be the most common podcast category in our dataset, with many of these shows consisting of recorded Christian sermons. Other broad topics appear to be further fragmented, such as Christianity, for which a number of topics appear to correspond to different parts of the liturgy or themes within Christianity.
By contrast, however, there are a number of topical communities that include creators from multiple categories or supersede category boundaries entirely. Two such communities focus on spirituality and self-improvement, which include content from episodes in the Religion, Society, and Business categories and the Business, Religion, and Health categories respectively. This is also the case for topics related to COVID-19, as well as racial justice, which is explored in-depth in §5. The presence of such topical communities suggests that existing category labels do not fully describe the meaningful groupings in podcast content, and that these are the areas in which we might expect to see the most extensive cross-fertilization and exchange of ideas.
@@
Guests play an important role in structuring parts of the podcast ecosystem. High-profile guests may appear frequently on multiple podcasts to promote their brand or ideology Bratcher and Cabosky (2024), facilitating the exchange or even diffusion of ideas. As a result, guest appearances are often a strong indicator of what type of content hosts prefer to feature and what will appeal to their audience.
We find that the association between guest appearances and category label varies significantly across categories, as shown by the podcast-guest network in Figure 2. Here, we visualize the network with nodes colored by category, positioned by the force-directed Yifan Hu algorithm, and sized proportionate to their edge degree Hu (2005). The figure shows that some categories such as Business, News, and Sports correspond to relatively distinct networked communities—i.e., podcasts within the same category share a distinct pool of guests that do not interact much with podcasts from different categories.
Despite being the largest categories, Religion and Society do not form large networked-communities in our network. As shown in supplemental Figure 9, both the Religion and Society categories appear to invite guests substantially less frequently than other large podcast categories, and thus provide fewer opportunities for cross-fertilization of ideas. This suggests that creators may be less well connected for these categories, though a longitudinal sample may help identify such structure.
@@
To quantify the relationship between category labels and identifiable networked communities, we calculate modularity for the partitions of our network corresponding to different categories to measure their intraconnectedness, as reported Table 1. This measures the extent to which membership in a category increases the probability of two nodes sharing a common guest relative to a graph with randomized edges (see Appendix H for details). The high modularity of Business and Sports suggests that podcasts within these categories are particularly likely to share guests relative to podcasts in other categories.
@@
Overall, we find evidence that categories provide meaningful structure to the podcast ecosystem. With respect to content, we find many distinct fandoms represented (e.g., Wrestling, Bitcoin), with many such topical communities belonging to a single category. However, a number of high-stakes topics (i.e. COVID-19, racial justice, and wellness) have cross-category discussion, a potential indicator that these domains may serve as a ground for discussion, debate, or the diffusion of ideas. With respect to guests, we find that Business and Sports form networked communities, whereas two of our largest categories (Religion, Society) do not. Our results cannot by themselves inform how community structure influences discourse in the podcast sphere. However, they answer the most fundamental question for future work by telling us where this community structure exists.
@@ 5 Collective Attention in the Ecosystem
In news domains, media environments often respond rapidly to major events, setting the agenda for public discussion and debate King et al. (2017); Boydstun et al. (2014). Due to their ubiquity, podcasts may play an equally important role in agenda-setting. However, given its fragmentation, it’s not currently known how responsive the ecosystem is to current events. Here, we test for collective attention using a case study on George Floyd’s murder, measuring both the (1) temporal responsiveness and (2) overall diffusion associated with this event. George Floyd’s death was one of the most significant events of that year in terms of news coverage Cowart et al. (2022); Reny and Newman (2021), and, as it occurs in the middle of the SPoRC time frame, provides an ideal test case of the response of the podcast ecosystem to real-world events.
@@ Results
Collective attention in podcasts responds rather quickly to the initial event, then decreases far slower over the following weeks (Figure 3). In particular, the percentages for both the George Floyd and Black Lives Matter topics rises to their peaks over the course of roughly 10 days, with a lag of approximately 4 days in-between them. This pattern of quick story development and slow story decay aligns with the phenomenon of a “media storm” in the news medium Boydstun et al. (2014). However, this case suggests that development of media storms in podcasts may operate at a slower pace, as prior computational work on news media found that the average storm had peaked by day three and decayed by day ten Litterer et al. (2023). Given that there are substantially more podcasts than media outlets (in part due to ease of entry), future work is needed to understand how the two ecosystems differ in their responses.
@@
Our analyses confirm that 1) the podcast ecosystem’s response had similar temporal patterns to news media, and that 2) discussion of George Floyd was widespread, yet variable, across communities. In the first case, our finding builds off of work on “media storms” and builds toward future work into the underlying mechanisms through which collective attention in podcasts rapidly shift and slowly decay. With regard to the widespread discussion of George Floyd, our work is broadly supported by findings that political discussion is abundant across communities with both political and non-political labels Rajadesingan et al. (2021); Munson and Resnick (2011). This suggests that, as with other platforms, listeners may be incidentally exposed to political content from hosts whom they trust and feel connected to Pew Research Center (2023b); Fletcher and Nielsen (2018); Weeks and Lane (2020).
@@ 6 Discussion
In addition to NLP applications, our data and analyses enable deeper investigation into social science questions. Section 4 emphasized structure, identifying communities within podcasts. Future research could investigate the political and social characteristics of these communities, similarly to past work on the Reddit and News ecosystems Waller and Anderson (2021); Niculae et al. (2015). For example, our textual and prosodic information could be used to infer whether communities form along the lines of gender performance and identity in this male-dominated space.
In section 5, we illustrate collective attention dynamics in the podcast ecosystem through our study of racial justice discussions in 2020. Rather than focusing on a single event, future work may consider how this ecosystem responds to a number of politically salient topics, extending work on news media to the podcast ecosystem Boydstun et al. (2014); Litterer et al. (2023). We found that discussion of George Floyd permeated even categories such as religion and business. Building on this finding, future work could model how information diffuses across communities and under what conditions this is likely to occur.
@@ 7 Conclusion
Podcasts are a widely distributed, unique, and impactful form of media, but comprehensive data for the computational study of this medium have been lacking. In this work, we present SPoRC, a dataset of over 1 million podcast transcripts alongside foundational analyses for the continued study of the podcast ecosystem. Our data is comprehensive, characterizing this medium in terms of its content, audio characteristics, speaker roles, and a number of other relevant factors such as category, hosting service, and publication date. Our structural analysis describes what content is discussed by which categories and demonstrates that podcast communities are formed through shared guests. Our temporal analysis finds that the ecosystem responded similarly to how the news media respond to major events. Furthermore, discussion of George Floyd was widespread, reaching a diverse set of categories and a large fraction of total podcasts. Our data and analyses answer important questions about podcasts. More importantly, however, they open the door to the continued research of this impactful and understudied medium.
@@
Our data and analysis are limited in a number of ways. With regard to our dataset, there are certain podcasts that are not publicly available and therefore excluded from our data collection pipeline. One noteworthy example is the popular Joe Rogan podcast, which was, until recently, available exclusively through Spotify. Moreover, despite being a large online repository, Podcast Index is not guaranteed to have comprehensive coverage of existing podcasts.
Our data processing pipeline involves additional limitations. Whisper models are predictive tools, meaning that they are inherently subject to mistakes and even hallucinations. These issues may be exacerbated in the context of non-majority accents, code-switching, or audio recorded with low-quality equipment or in noisy environments. In order to maximize data quality, we remove cases of potential hallucination, but this may disproportionately affect already marginalized groups. Furthermore, our diarization and role annotation approaches are subject to error that is difficult to quantify and could propagate to downstream analyses in unforeseen ways. In addition, some hosts do not introduce themselves, hence we are unable to identify them, and we do not attempt to resolve names to speakers in the case of more than one host or guest.
While we provide a first large-scale analysis of podcasts, our conclusions are limited. In our analyses, we use a subset of data that is dense but represents a narrow slice of time. Thus, our conclusions may not generalize outside this timeframe. Further limitations with regard to our temporal analysis stem from our inability to pinpoint mechanistic factors underlying our observed time series. While prior work is suggestive of particular explanations for the patterns observed within and across categories, our data do not permit a definitive answer for “why” these patterns emerge.
@@
At the data processing phase, one potential harm arises if a speaker’s actual speech is mistranscribed or misattributed. Such a mistake would be particularly problematic if this speech is offensive or reflects stereotypes or other biases. Furthermore, this type of error could be more likely for users with non-majority ways of speaking, which could exacerbate this harm for groups that are already marginalized. To mitigate this issue at the transcription level, we have undertaken thorough manual review to ensure that hallucinations are rare and that repeated text is filtered from our data. In the case of diarization, this issue is more difficult to mitigate. Recognizing that the potential for harm due to speaker misattribution varies widely across contexts, we encourage future users of our data to consider how this potential error could manifest when interpreting downstream results.
We further identify two potential harms due to the open release of our data. First, releasing our data could result in privacy concerns. Although we only process podcasts that are publicly available, the release of our processed data could expose a podcast’s content to an unforeseen or undesired audience. Creators may have expectations that their content will not be widely distributed, and the release of our data could violate this expectation. Furthermore, the enhanced searchability afforded by podcast transcription could render a podcast’s content easier to access than was originally intended. A second and related harm is the usage of data for training of large language models. Some creators may not want their content being used to train deep learning systems, and the release of our data makes content more accessible for these purposes. This issue is particularly concerning when content is used to generate revenue without compensating the original creators.
@@
Acknowledgments
This work was supported in part by the National Science Foundation under Grant No. IIS-2143529 and by DSO National Laboratories.
Appendix